# The Baldwin Effect in Advancing Generalizability of Physics-Informed Neural Networks

### v2

@@ Abstract
Physics-informed neural networks (PINNs) are at the forefront of scientific machine learning, making possible the creation of machine intelligence that is cognizant of physical laws and able to accurately simulate them. However, today’s PINNs are often trained for a single physics task and require computationally expensive re-training for each new task, even for tasks from similar physics domains. To address this limitation, this paper proposes a pioneering approach to advance the generalizability of PINNs through the framework of Baldwinian evolution. Drawing inspiration from the neurodevelopment of precocial species that have evolved to learn, predict and react quickly to their environment, we envision PINNs that are pre-wired with connection strengths inducing strong biases towards efficient learning of physics. A novel two-stage stochastic programming formulation coupling evolutionary selection pressure (based on proficiency over a distribution of physics tasks) with lifetime learning (to specialize on a sampled subset of those tasks) is proposed to instantiate the Baldwin effect. The evolved Baldwinian-PINNs demonstrate fast and physics-compliant prediction capabilities across a range of empirically challenging problem instances with more than an order of magnitude improvement in prediction accuracy at a fraction of the computation cost compared to state-of-the-art gradient-based meta-learning methods. For example, when solving the diffusion-reaction equation, a 70x improvement in accuracy was obtained while taking 700x less computational time. This paper thus marks a leap forward in the meta-learning of PINNs as generalizable physics solvers. Sample codes are available at https://github.com/chiuph/Baldwinian-PINN.
@@
Baldwin effect, neuroevolution, meta-learning, physics-informed neural networks
@@ II-A2 Generalizable neural physics solver
While most PINN models are trained to solve a specific physics task, there is increasing interest in generalizable neural physics solvers, i.e. models that can be flexibly applied to multiple physics problems once trained. In this context, we can consider a physical phenomena of interest that is represented by a set of training tasks belonging to some underlying task-distribution $p(\bm{\mathcal{T}})$, e.g., a family of PDEs spanning different PDE parameters, ICs and/or BCs.
@@ III Proposed Methodology
We address the above limitations with a novel meta-learning PINN framework based on Baldwinian neuroevolution. Baldwinian neuroevolution aims to optimize the weight distribution in the pre-final nonlinear hidden layers and physics-informed learning hyperparameters that are essential for achieving optimal performance in task-specific physics-informed learning. We limit our task-specific learning to the output layer, allowing us to solve linear problems with a 1-step pseudoinverse operation (or a few steps for nonlinear problems) during both the meta-learning process and test time, in a consistent manner. Importantly, once given the nonlinear hidden layers’ weights and learning hyperparameters, the pseudoinverse operation is exact and fast (does not require further gradient descent updates). An illustration of Baldwinian neuroevolution for PINN is shown in Figure 2(a).
III-A A two-stage stochastic programming formulation
@@
where ${\bf\it w}=[\dots w_{\_}j\dots]^{T}$ are the output layer weights, and $f_{\_}j(x,t;{\bf\it\tilde{w}})$’s represent nonlinear projections of the input variables with the hidden layers’ weights ${\bf\it\tilde{w}}$. The connection strengths ${\bf\it\tilde{w}}$ are typically learned during PINN training. However, in the proposed implementation of Baldwinian-PINN as detailed below, they are drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$ (see example in Section III-B), that are assigned at birth as per Baldwinian evolution.
The model’s reaction to a given environment—i.e., finding the best set of $w_{\_}j$’s such that the Baldwinian-PINN’s output satisfies Eq. 1 for a specific physics task $\mathcal{T}_{\_}i$—is the focus of a typical single PINN problem described in Section II-A. This is reduced to a physics-based least-squares problem in this work:
$\operatorname*{arg\,min}_{\_}{{\bf\it w}}\ \mathbf{(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})^{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})+\lambda{\bf\it w}^{\mathbf{T}}{\bf\it w}$ (3)
where $\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}$ is obtained by substituting the model’s output into the left hand side of Eq. 1 for a given set of collocation points, and $\mathbf{b}^{\mathcal{T}_{\_}i}$ represents the corresponding right hand side of Eq. 1. This is chosen as Eq. 3 yields a closed-form result when using the Moore-Penrose generalized inverse, permitting extremely fast learning (order of milli-seconds in our experiments) for physics-compliant prediction. Similar least-squares formulations studied in the literature have been shown to be competitive with, or faster than, widely-used numerical solvers such as the finite element method [34, 35]. The learning hyperparameter $\lambda\geq 0$ reduces the $L^{2}$-norm of the least-squares solution, thereby improving the solution numerically. Detailed derivations, including information on the construction of matrix $\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}$ and vector $\mathbf{b}^{\mathcal{T}_{\_}i}$, are provided in Section III-C.
@@
where the solution ${\bf\it w^{*}}$ of the second stage problem:
${\bf\it w^{*}}=\operatorname*{arg\,min}_{\_}{{\bf\it w}}\ \mathbf{(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})^{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})+\lambda{\bf\it w}^{\mathbf{T}}{\bf\it w}$ (5)
provides optimal output layer weights that allow the model to specialize to any realization of task $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}})$ for the given network’s ${\bf\it\tilde{w}}\backsim p_{\_}{\theta}({\bf\it\tilde{w}})$.
@@
$l_{\_}{LSE}({\bf\it w^{*}})=\mathbf{(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w^{*}}-\mathbf{b}^{\mathcal{T}_{\_}i})^{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w^{*}}-\mathbf{b}^{\mathcal{T}_{\_}i})$ (6)
@@
III-B Baldwinian-PINN instantiation: shallow network architecture and connections at birth
@@
In the present study, Baldwinian-PINNs are designed to have weights and biases in the nonlinear hidden layers fixed to random values at birth—a randomized neural networks setup [36, 37]—drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$. Both normal and uniform distributions for the weights and biases are possible for the randomized PINNs [34, 35]. Similarly, smooth nonlinear activations such as sin, softplus, and tanh are common in PINN literature, and can be advantageous for different problems. For greater flexibility, we apply both distributions for setting weights and biases and all three activations to different hidden layer blocks in the Baldwinian-PINN models.
@@
Hence, the base Baldwinian-PINNs use a single hidden layer architecture which is segmented into $3\times 2=6$ unique blocks, with the weights and biases from 3 blocks sampling from normal distributions and the weights and biases from another 3 blocks sampling from uniform distributions. Each block has a fixed number of neurons ($n_{\_}{neuron}$ = 150 or 200). Assuming 2 input variables $(x,t)$, we can write the output $f_{\_}j$ for all the neurons $j=1,...,n_{\_}{neuron}$ in a nonlinear hidden-layer block as:
@@
The base Baldwinian-PINN model is depicted in Figure 2(b). Given the Baldwinian-PINNs’ configuration, a set of network hyperparameters $\theta=(m^{1},s^{1},...,m^{18},s^{18})$ control the distributional mean and spread of the weights and biases in different blocks. In the spirit of the Baldwin effect, the task-specific output layer ${\bf\it w}$ is the outcome of lifetime learning and not directly inherited by the next generation of offspring [38]; only the hyperparameters $(\theta,\lambda)$ are subjected to evolutionary variation and inheritance. Since the neuroevolution only searches the distribution parameters for groups of weights instead of evolving each individual weight in the nonlinear hidden layer, the reduced dimensionality can be even more effectively searched by today’s evolutionary algorithms.
@@
III-C Baldwinian-PINN lifetime learning
$\mathbf{A}{\bf\it w}$ $=\mathbf{b}$ (36)
${\bf\it w}$ $=(\lambda I+\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}\mathbf{b}$ (37a)
(if the system of equations in Eq. 36 is over-determined)
${\bf\it w}$ $=\mathbf{A}^{T}(\lambda I+\mathbf{A}\mathbf{A}^{T})^{-1}\mathbf{b}$ (37b)
(if the system of equations in Eq. 36 is under-determined)
For nonlinear PDEs, iterative methods can be used for arriving at optimized $w_{\_}j$’s in Baldwinian-PINNs. In this work, we use a lagging of coefficient approach which is common in numerical methods [40] and detailed in Suppl. LABEL:subsec:lagging. Briefly, we approximately linearize the nonlinear term(s) in Eq. 1 by substituting the output $u(x,t)=\sum_{\_}jw_{\_}jf_{\_}j(x,t;{\bf\it\tilde{w}})$ obtained from previous step, and iteratively solve Eq. 37 to update $w_{\_}j$’s for a fixed number of steps, $N$, or until a convergence criterion is reached.
Hence, the Baldwinian-PINNs’ lifetime learning is broadly applicable to both linear and nonlinear PDEs, with the psuedoinverse formulation permitting extremely fast computation for physics-compliant prediction. In addition, there is no requirement on prior labelled data for any new task as the learning can be entirely physics-based (as per conventional PINNs).
III-D Baldwinian neuroevolution algorithm and hyperparameters
Algorithm 1 and Algorithm 2 describe a Baldwinian neuroevolution procedure to solve the two-stage stochastic programming problem defined in Section III-A. Recall a Baldwinian-PINN individual can be represented by $(\theta,\lambda)$, i.e., network architecture and learning algorithm hyperparameters, during neuroevolution.
Algorithm 1 Baldwinian neuroevolution of PINNs
1:$p(\bm{\mathcal{T}})$: distribution over physics tasks
2:$\bm{\mathcal{P}}$: initial population-representation of individuals (i.e., network architecture and learning algorithm hyperparameters in Baldwinian-PINN models)
3:$\bm{\mathcal{S}}$: procedure to obtain a batch of individuals given a population-representation
4:$\bm{\mathcal{U}}$: procedure to update a population-representation given a batch of fitness-scored individuals
$\bm{\mathcal{F}}$: procedure to return fitness based on lifetime learning performance of individual given a set of tasks
6:while not done do
7:     Sample batch of individuals from population $(\theta^{g},\lambda^{g})\backsim\bm{\mathcal{S}}(\bm{\mathcal{P}}),g=1,...,n_{\_}{pop}$
8:     Sample batch of tasks $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}}),i=1,...,n_{\_}{task}$
9:     for all $(\theta^{g},\lambda^{g})$ do
[…]
11:     end for
12:     Update population based on fitness of individuals:
13:      $\bm{\mathcal{P}}\leftarrow\bm{\mathcal{U}}(\bm{\mathcal{P}},\{(\theta^{g},\lambda^{g},f^{g}),g=1,....,n_{\_}{pop}\})$
14:end while
INPUT: network architecture and learning algorithm hyperparameters $(\theta,\lambda)$, list of tasks $\mathcal{T}_{\_}i,i=1,...,n_{\_}{task}$
OUTPUT: fitness $f$
1:$\bm{\mathcal{G}}$: procedure to sample weights and biases in nonlinear hidden-layer(s)
2:$\bm{\mathcal{C}}$: procedure to construct least squares problem on a set of collocation points based on underlying physics (PDEs, BCs, ICs) of the task
3:$N$: number of iterations to take during per-task “lifetime learning”: $1$ for linear PDEs, $>1$ for nonlinear PDEs
4:Sample hidden layers’ weights and biases ${\bf\it\tilde{w}}=\bm{\mathcal{G}}(\theta)$
5:for all $\mathcal{T}_{\_}i$ do
6:     ${\bf\it w}=[\dots w_{\_}j\dots]^{T}=\mathbf{0}$triangleright initial solution for nonlinear PDEs
7:     for $k=1,...,N$ do
8:         Construct least squares matrix and vector $(\mathbf{A},\mathbf{b})=\bm{\mathcal{C}}(\mathcal{T}_{\_}i,{\bf\it\tilde{w}},{\bf\it w})$
9:         Compute least-squares solution via pseudoinverse:
10:            ${\bf\it w}=(\lambda I+\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}\mathbf{b}$ ,
11:            if $\mathbf{A}$ has more rows than columns
12:            ${\bf\it w}=\mathbf{A}^{T}(\lambda I+\mathbf{A}\mathbf{A}^{T})^{-1}\mathbf{b}$ ,
13:            if $\mathbf{A}$ has more columns than rows
14:     end for
15:     ${\bf\it w}^{\mathcal{T}_{\_}i}\leftarrow{\bf\it w}$, $\mathbf{A}^{\mathcal{T}_{\_}i}\leftarrow\mathbf{A}$, $\mathbf{b}^{\mathcal{T}_{\_}i}\leftarrow\mathbf{b}$
16:end for
[…]
18:Compute mean squared error: $l_{\_}{MSE}=\sum_{\_}{\mathcal{T}_{\_}i}\frac{1}{n}\sum_{\_}s^{n}(u_{\_}s^{label}-\sum_{\_}jw_{\_}j^{\mathcal{T}_{\_}i}f_{\_}j(x_{\_}s,t_{\_}s;{\bf\it\tilde{w}}_{\_}j))^{2}$ over the labelled data $u_{\_}s^{label},s=1,...,n$
19:Compute overall fitness: $f=-(\tau_{\_}{LSE}\ l_{\_}{LSE}+\tau_{\_}{MSE}\ l_{\_}{MSE})$
The Baldwinian neuroevolution procedure described in Algorithm 1 is generic for neuroevolution methods, such as covariance matrix adaptation evolution strategy (CMA-ES) [41]. As an instantiation of information-geometric optimization algorithms [42], CMA-ES represents the population of $(\theta,\lambda)$ in $\bm{\mathcal{P}}$ using a multivariate normal search distribution, initialized with zeros mean and standard deviation (std.) as tuning hyperparameter. $\bm{\mathcal{S}}$ becomes a random sampling procedure with population size or the number of individuals $n_{\_}{pop}$ to be sampled from the search distribution as hyperparameter, and $\bm{\mathcal{U}}$ is the CMA-ES update-step [41] which iteratively adapts its search distribution based on the rank-based fitness landscape until the convergence criteria, e.g., a pre-determined number of iterations or fitness value (trade-off between computation resource and convergence) is met.
The network weights and biases can take any value from (($\infty$( - ∞, $\infty$)∞ ), hence there is no restriction to the continuous search space of $\theta=(m^{1},s^{1},...)$ representing their distributional means and spreads. The learning hyperparameter $\lambda\geq 0$ can be evolved in continuous search space and its absolute value is then used for computing the least-squares solution. In our implementation with CMA-ES, $(\theta,\lambda)$ share the same initial standard deviation, and we scale the learning hyperparameter by a factor, i.e., $\lambda\leftarrow 1\mathrm{e}{-4}\times\text{abs}(\lambda)$, to improve the performance of the pseudoinverse.

### v3

# Evolutionary Optimization of Physics-Informed Neural Networks: Advancing Generalizability by the Baldwin Effect

@@ Abstract
Physics-informed neural networks (PINNs) are at the forefront of scientific machine learning, making possible the creation of machine intelligence that is cognizant of physical laws and able to accurately simulate them. However, today’s PINNs are often trained for a single physics task and require computationally expensive re-training for each new task, even for tasks from similar physics domains. To address this limitation, this paper proposes a pioneering approach to advance the generalizability of PINNs through the framework of Baldwinian evolution. Drawing inspiration from the neurodevelopment of precocial species that have evolved to learn, predict and react quickly to their environment, we envision PINNs that are pre-wired with connection strengths inducing strong biases towards efficient learning of physics. A novel two-stage stochastic programming formulation coupling evolutionary selection pressure (based on proficiency over a distribution of physics tasks) with lifetime learning (to specialize on a sampled subset of those tasks) is proposed to instantiate the Baldwin effect. The evolved Baldwinian-PINNs demonstrate fast and physics-compliant prediction capabilities across a range of empirically challenging problem instances with more than an order of magnitude improvement in prediction accuracy at a fraction of the computation cost compared to state-of-the-art gradient-based meta-learning methods. For example, when solving the diffusion-reaction equation, a 70x improvement in accuracy was obtained while taking 700x less computational time. This paper thus marks a leap forward in the evolutionary meta-learning of PINNs as generalizable physics solvers. Sample codes are available at https://github.com/chiuph/Baldwinian-PINN.
@@
Baldwin effect, evolutionary optimization, neuroevolution, meta-learning, physics-informed neural networks
@@ II-A2 Generalizable neural physics solver
While most PINN models are trained to solve a specific physics task, there is increasing interest in generalizable neural physics solvers, i.e. models that can be flexibly applied to multiple physics problems once trained. In this context, we can consider a physical phenomena of interest that is represented by a set of training tasks belonging to some underlying task-distribution $p(\bm{\mathcal{T}})$, e.g., a family of PDEs spanning different PDE parameters $\vartheta$, different ICs $u_{\_}0(x)$ and/or different BCs $g(x,t)$.
@@ III Proposed Methodology
We address the above limitations with a meta-learning PINN framework based on Baldwinian neuroevolution. A novel evolutionary algorithm is crafted to jointly optimize the weight distribution in the network’s hidden layers alongside other hyperparameters that are essential for achieving optimal performance on downstream physics-informed learning tasks. We limit our task-specific learning to the output layer, allowing us to solve linear problems with a 1-step pseudoinverse operation (or a few steps for nonlinear problems) during both the meta-learning process and test time, in a consistent manner. Importantly, once given the nonlinear hidden layers’ weights and learning hyperparameters, the pseudoinverse operation is exact and fast (does not require further gradient descent updates). An illustration of the Baldwinian evolution of PINNs is shown in Figure 2(a).
III-A The Baldwin effect as a two-stage stochastic program
@@
where ${\bf\it w}=[\dots w_{\_}j\dots]^{T}$ are the output layer weights, and $f_{\_}j(x,t;{\bf\it\tilde{w}})$’s represent nonlinear projections of the input variables with the hidden layers’ weights ${\bf\it\tilde{w}}$. The connection strengths ${\bf\it\tilde{w}}$ are typically learned during PINN training. However, in the proposed implementation of Baldwinian-PINNs as detailed below, they are drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$—see example in Section III-D1 and Figure. 2(b)—that are assigned at birth as per Baldwinian evolution.
The model’s reaction to a given environment—i.e., finding the best set of $w_{\_}j$’s such that the model’s output satisfies Eq. 1 for a specific physics task $\mathcal{T}_{\_}i$—is the focus of a typical PINN described in Section II-A. This is reduced to a physics-based least-squares problem in this work:
${\bf\it w^{*}}=\operatorname*{arg\,min}_{\_}{{\bf\it w}}\ (\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})^{\mathbf{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})+\lambda{\bf\it w}^{\mathbf{T}}{\bf\it w}$ (3)
where $\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}$ is obtained by substituting the model’s output into the left hand side of Eq. 1 for a given set of collocation points, and $\mathbf{b}^{\mathcal{T}_{\_}i}$ represents the corresponding right hand side of Eq. 1. Eq. 3 yields a closed-form solution ${\bf\it w^{*}}$ when using the Moore-Penrose generalized inverse, permitting extremely fast learning (order of milli-seconds in our experiments) for physics-compliant prediction. Similar least-squares formulations studied in the literature have been shown to be competitive with, or faster than, widely-used numerical solvers such as the finite element method [34, 35]. The learning hyperparameter $\lambda\geq 0$ reduces the $L^{2}$-norm of the least-squares solution, thereby improving the solution numerically. Detailed derivations, including information on the construction of matrix $\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}$ and vector $\mathbf{b}^{\mathcal{T}_{\_}i}$, are provided in Section III-C.
@@
where ${\bf\it w^{*}}$ is the solution of the second stage problem defined earlier in Eq. 3, which provides optimal output layer weights that allow the model to specialize to any realization of task $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}})$ for the given network’s ${\bf\it\tilde{w}}\backsim p_{\_}{\theta}({\bf\it\tilde{w}})$.
@@
$l_{\_}{LSE}({\bf\it w^{*}})=(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w^{*}}-\mathbf{b}^{\mathcal{T}_{\_}i})^{\mathbf{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w^{*}}-\mathbf{b}^{\mathcal{T}_{\_}i})$ (5)
@@
In what follows, we present a novel procedure for solving the two-stage stochastic program via an evolutionary algorithm. The hyperparameters $\theta,\lambda$ are evolved from one generation to the next, with their fitness evaluations for the first stage problem conditioned on lifetime learning to obtain the optimal ${\bf\it w^{*}}$ from the second stage problem. Note that while the outcome of lifetime learning influences the selection pressure acting on the evolving hyperparameters, it does not directly alter the genetic makeup of the hyperparameters within a generation. As such, our overall method exhibits a clear connection with the evolutionary principles of Baldwinism.
III-B Baldwinian neuroevolution
The Baldwinian neuroevolution procedure to solve the two-stage stochastic programming problem defined in Section III-A is described in Algorithm 1 and Algorithm 2. Recall from Figure. 2(b) that a Baldwinian-PINN is represented by $(\theta,\lambda)$, i.e., distribution of network weights and biases and lifetime learning hyperparameters, during neuroevolution.
Algorithm 1 Baldwinian neuroevolution of PINNs
INPUT: training tasks distribution, $p(\bm{\mathcal{T}})$
OUTPUT: best solution found (i.e., Baldwinian-PINN model’s weights and biases in nonlinear hidden-layers and lifetime learning hyperparameters $(\theta,\lambda)$)
$\bm{\mathcal{F}}$: procedure to return fitness based on lifetime learning performance of individual given a batch of tasks (described in detail in Algorithm 2)
2:Initialize population $\bm{\mathcal{P}}$
3:while not done do
4:     Sample a new batch of offspring $(\theta^{g},\lambda^{g}),g=1,...,n_{\_}{pop}$ from the population $\bm{\mathcal{P}}$
5:     Sample batch of tasks $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}}),i=1,...,n_{\_}{task}$
6:     for all $(\theta^{g},\lambda^{g})$ do
[…]
8:     end for
9:     Update population $\bm{\mathcal{P}}$ based on fitness of individuals: $\{(\theta^{g},\lambda^{g},f^{g}),g=1,....,n_{\_}{pop}\}$
10:end while
11:Return best individual $(\theta,\lambda)$ found in population $\bm{\mathcal{P}}$
INPUT: network and lifetime learning hyperparameters $(\theta,\lambda)$, batch of tasks $\mathcal{T}_{\_}i,i=1,...,n_{\_}{task}$
OUTPUT: fitness $f$
1:$\bm{\mathcal{G}}$: procedure to sample weights and biases in nonlinear hidden-layers (exemplified in Section III-D1 and Figure. 2(b))
2:$\bm{\mathcal{C}}$: procedure to construct least squares problem on a set of collocation points based on underlying physics (PDEs, BCs, ICs) of the task (described in detail in Section III-C)
3:Sample hidden layers’ weights and biases ${\bf\it\tilde{w}}=\bm{\mathcal{G}}(\theta)$
4:for all $\mathcal{T}_{\_}i$ do
5:     Construct least squares matrix and vector:
6:     $(\mathbf{A},\mathbf{b})=\bm{\mathcal{C}}(\mathcal{T}_{\_}i,{\bf\it\tilde{w}})$
7:     Compute least squares solution as per Eq. 34:
8:     ${\bf\it w^{*}}=(\lambda I+\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}\mathbf{b}$
9:     Compute least squares error (LSE) as per Eq. 5:
10:     $l_{\_}{LSE}^{\mathcal{T}_{\_}i}=(\mathbf{A}{\bf\it w^{*}}-\mathbf{b})^{\mathbf{T}}(\mathbf{A}{\bf\it w^{*}}-\mathbf{b})$
11:     Compute mean squared error (MSE) as per Eq. 6:
12:     $l_{\_}{MSE}^{\mathcal{T}_{\_}i}=\frac{1}{n}\sum_{\_}s^{n}(u_{\_}s^{label}-\sum_{\_}jw_{\_}j^{*}f_{\_}j(x_{\_}s,t_{\_}s;{\bf\it\tilde{w}}_{\_}j))^{2}$
13:end for
14:Compute overall fitness: $f=-(\tau_{\_}{LSE}\ \sum_{\_}{\mathcal{T}_{\_}i}l_{\_}{LSE}^{\mathcal{T}_{\_}i}+\tau_{\_}{MSE}\ \sum_{\_}{\mathcal{T}_{\_}i}l_{\_}{MSE}^{\mathcal{T}_{\_}i})$
The Baldwinian neuroevolution procedure described in Algorithm 1 is generic for evolutionary optimization methods. The algorithm initializes a population $\bm{\mathcal{P}}$ of Baldwinian-PINN models given by different weights and biases in the nonlinear hidden-layers and lifetime learning hyperparameters $(\theta,\lambda)$. For a probabilistic model-based evolutionary algorithm, $\bm{\mathcal{P}}$ is commonly represented by a search distribution. In each generation, $n_{\_}{pop}$ new offspring individuals are sampled from the search distribution (probabilistic model-based EAs) or through crossover/mutation (traditional EAs), and their fitness are evaluated for a batch of tasks randomly sampled from the training task distribution $p(\bm{\mathcal{T}})$. The fitness evaluation procedure $\bm{\mathcal{F}}$ gives the lifetime learning outcome $f^{g}$ of these offspring individuals $(\theta^{g},\lambda^{g}),g=1,...,n_{\_}{pop}$ for the given tasks. In line with the essence of Baldwinism, the lifetime learning procedure (described in detail in Algorithm 2) does not alter the genetic makeup $(\theta,\lambda)$ of the individuals. The outcome of the lifetime learning procedure specifies the fitness which creates the selection pressure influencing the evolution of the population, but is not directly inherited by the next population (unlike Lamarckian evolution). The Baldwinian neuroevolution algorithm iteratively adapts $\bm{\mathcal{P}}$ towards offspring with better fitness until the convergence criteria, e.g., a pre-determined number of generations or fitness value (trade-off between computation resource and convergence), is met.
Given an individual $(\theta,\lambda)$ sampled from the search distribution and a batch of tasks sampled from the training distribution, $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}}),i=1,...,n_{\_}{task}$, the procedure $\bm{\mathcal{F}}$ to return fitness is detailed in Algorithm 2. It starts with a procedure $\bm{\mathcal{G}}$ to populate nonlinear hidden-layers’ weights and biases ${\bf\it\tilde{w}}$ of a Baldwinian-PINN from the sampled individual $\theta$. In the present study, Baldwinian-PINNs are designed to have ${\bf\it\tilde{w}}$ fixed at birth to random values drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$, akin to a randomized neural networks setup [36, 19]. This procedure is exemplified in Section III-D1. Then, lifetime learning of the Baldwinian-PINN is performed to obtain the optimal network weights ${\bf\it w^{*}}$ in the linear output layer, for each of the sampled tasks $\mathcal{T}_{\_}i$. It involves procedure $\bm{\mathcal{C}}$ to construct least squares problem, i.e., matrix and vector $(\mathbf{A},\mathbf{b})$, on a fixed set of collocation points based on underlying physics (PDEs, BCs, ICs) of the task. This procedure is detailed in Section III-C. We choose the collocation points to coincide with the location of labelled data for MSE computation, although this is not a prerequisite. Finally, the overall fitness $f$ can be computed by aggregating the LSE and MSE based on lifetime learning outcomes for all the sampled tasks.
In the spirit of the Baldwin effect, the task-specific output layer ${\bf\it w^{*}}$ is the outcome of lifetime learning and not directly inherited by the next generation of offspring [37]; only the hyperparameters $(\theta,\lambda)$ are subjected to evolutionary variation and inheritance.
III-C Baldwinian-PINN lifetime learning procedure
$\mathbf{A}{\bf\it w}$ $=\mathbf{b}$ (33)
${\bf\it w^{*}}=(\lambda I+\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}\mathbf{b}$ (34)
For nonlinear PDEs, iterative methods can be used for arriving at optimized $w_{\_}j$’s in Baldwinian-PINNs (detailed in Suppl. LABEL:subsec:lagging). Hence, the Baldwinian-PINNs’ lifetime learning is broadly applicable to both linear and nonlinear PDEs, with the psuedoinverse formulation permitting extremely fast computation for physics-compliant prediction. It is worth emphasizing that each Baldwinian-PINN undergoes lifetime learning in order to produce accurate, physics-compliant predictions on a single, new physics task, hence, the fast nature of this procedure permits flexible and rapid prediction for each new task on-demand. In addition, there is no requirement on prior labelled data for any new task as the learning can be entirely physics-based (as per conventional PINNs).
III-D Implementation details
III-D1 Baldwinian-PINN architecture
@@
In the present study, Baldwinian-PINNs are designed to have weights and biases in the nonlinear hidden-layers ${\bf\it\tilde{w}}$ fixed to random values at birth. They are akin to a randomized neural networks setup [36, 19] and drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$. Both normal and uniform distributions for the weights and biases are possible for the randomized PINNs [34, 35]. Similarly, smooth nonlinear activations such as sin, softplus, and tanh are common in PINN literature, and can be advantageous for different problems. For greater flexibility, we apply both distributions for setting weights and biases and all three activations to different hidden layer blocks in the Baldwinian-PINN models.
@@
Our base Baldwinian-PINN model is depicted in Figure 2(b). It’s hidden layer architecture is segmented into $3\times 2=6$ unique blocks, with the weights and biases of the first 3 blocks sampled from normal distributions and the weights and biases of the other 3 blocks sampled from uniform distributions. Each block has a fixed number of neurons ($n_{\_}{neuron}$ = 150 or 200). Assuming 2 input variables $(x,t)$, we can write the output $f_{\_}j$ for all the neurons $j=1,...,n_{\_}{neuron}$ in a nonlinear hidden-layer block as:
@@
Given the Baldwinian-PINNs’ configuration, a set of network hyperparameters $\theta=(m^{1},s^{1},...,m^{18},s^{18})$ control the distributional mean and spread of the weights and biases in different blocks. Since the neuroevolution only searches the distribution parameters for groups of weights instead of evolving each individual weight in the nonlinear hidden layer, the reduced dimensionality can be even more effectively searched by today’s evolutionary algorithms.
@@
III-D2 Evolutionary algorithm